Evaluate


lim 3nΣr=1 n^2/(4n+r)^3


n→∞



1
Expert's answer
2021-12-14T15:01:04-0500

By integral test for series,

"lim_{n\to\infty} \Sigma^{3n}_1 \frac{n^2}{(4n+x)^3}\\
=lim_{n\to\infty} \Sigma^{3n}_1 \frac{1}{n(4+x)^3}\\
=\int^3_1 (x+4)^{-3}dx\\
=\frac{1}{-2}[(x+4)^{-2}]_1^3\\
=\frac{1}{-2}[\frac{1}{49}-\frac{1}{25}]\\
=\frac{2}{1225}"